140811 LML Observations

Isaac arrived at the LML site at 9:40am. The doors to the collector had been deployed and there was a small sample present but not enough for further analysis. The small sample was discarded. The collector was then disassembled and washed with 409 followed by DI water. A significant amount of bugs had accumulated on the strand frames which were removed during the washing process. The body of the collector was then washed and rinsed with DI. The collector was then reset for the next fog event.

Leave a Reply

Your email address will not be published. Required fields are marked *